# Quantum information with quantum-like bits

### v4

@@ Abstract
In previous work we have proposed a construction of quantum-like bits that could endow a large, complex classical system, for example of oscillators, with quantum-like function that is not compromised by decoherence. In the present paper we investigate further this platform of quantum-like states. Firstly, we discuss a general protocol on how to construct synchronizing networks that allow for emergent states. We then study how gates can be implemented on those states. This suggests the possibility of quantum-like computing on specially-constructed classical networks. Finally, we define a notion of measurement that allows for non-Kolmogorov interference, a feature that separates our model from a classical probabilistic system. This paper aims to explore the mathematical structure of quantum-like resources, and shows how arbitrary gates can be implemented by manipulating emergent states in those systems.
@@ I Introduction
In the spirit of quantum-inspired computing, we showed in recent work that a large class of synchronizing classical systems can be utilized as a resource for information processing. Scholes (2020, 2023, 2024); Scholes and Amati (2024) We focused, in particular, on many body systems whose correlations obey the graph structure shown in panel a) of ??. Here, two densely connected regions (in blue) of a network are mutually interacting via sparse links (in red). A special feature of this class of networks is that they produce pairs of emergent states, obtained by the direct sum of the emergent states of the blue subgraphs, with identical and opposite phases. The key requirement of the blue subgraphs, then, is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs. Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) In our framework, emergent states can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??). Their robustness against noise and stochastic disorder increases with system size, exhibiting a favorable trend in contrast to the typical behavior of genuine quantum resources. Scholes (2023, 2020) This intriguing feature is connected to the occurrence of classical synchronization phenomena. Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) We showed in Ref. Scholes, 2024 that the emergent states of connected expander subgraphs are isomorphic to the Hilbert space of a qubit. This led us to introducing the notion of quantum-like (QL) bits within the context of classically-synchronizing networks. In Ref. Scholes and Amati, 2024 we built on these results and proposed a general framework to construct resources for multiple QL bits. In particular, we showed that the Cartesian product ($\square$) of ${N_{\mathrm{QL}}}$ bipartite graphs obeys the same dimensionality of a register of ${N_{\mathrm{QL}}}$ qubits [see panel b) of ??].
In the present paper, we build on the foundational work of Ref. Scholes, 2024; Scholes and Amati, 2024; Scholes, 2023 to formalize the concept of QL states and gates. We present a computational framework for implementing information processing within the QL formalism through appropriate manipulations of the emergent states of network resources. As depicted in panel c) of ??, we show that an arbitrary gate transformation $V_{\mathrm{g}}$ can be expressed both as graph transformation, and as a map of the emergent states of the graph. Finally, we propose a measurement approach upon information processing that aligns with the QL probabilistic framework developed by Khrennikov Khrennikov (2007a, 2019, 2010). This theory shows that genuine quantum systems are not strictly required to process information in a QL way. Instead, to observe QL phenomena it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and their measurement. Consistently with Khrennikov’s approach, Khrennikov (2010) we list a set of measurement axioms which allows to define a generalized notion of nonclassical interference between observable operators.
@@
The present paper is structured as follows. In ?? we summarize the formalism describing how to construct a single-QL-bit Hilbert space. In ?? we discuss its generalization to higher dimensions, and we show explicitly in ?? how to construct Bell states for two QL bits. In ?? we study the action of a complete set of QL gates, and we provide concrete examples of their action. In ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach. In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models. Khrennikov (2010) Finally, in ??, we outline the complete workflow for QL information processing and discuss its efficiency and accuracy.
@@
In conjunction with Refs. Scholes, 2023; Scholes and Amati, 2024; Scholes, 2024, this work provides the mathematical foundation for a rigorous approach to QL information processing from classical synchronizing systems.
@@
Figure 1: Illustration of the main concepts presented in this paper and in prior work by the authors. Scholes (2024); Scholes and Amati (2024) a) Many body systems consisting of networks of synchronizing dipoles can easily exhibit classical synchronization phenomena. By structuring two-body correlations into two densely connected regions (blue) weakly interacting through sparse links (red), synchronization gives rise to pairs of emergent states. These states are well-distinguishable from a large band of random states at lower energies (dashed line in the sketch of the spectrum). In Ref. Scholes, 2024 we proposed a map from emergent states to the computational basis of a qubit. b) In Ref. Scholes and Amati, 2024, we demonstrated that the emergent states of the Cartesian product of a set of synchronizing graphs scale with the same dimensionality of a qubit system. In this paper, we expand over these concepts by providing further mathematical details on the formalism for this many-body problem. c) Additionally, we elaborate on the concept of information processing within the QL framework. We show that operations on graphs, implemented through suitable transformations of their edges, are equivalent to gate transformations $V_{\mathrm{g}}$ applied directly to the emergent states.
@@ II Quantum-like bits
In this section we derive a mapping from a classical network to the Hilbert space of interacting QL bits. While the concept has been already elucidated in Ref. Scholes, 2024, here we lay out a more formal description. Later in the paper we use this formalism to study the notion of gates to perform quantum information processing, and we provide concrete examples.
@@ II.1 Isolated system
Let us consider two classical many-body systems with pairwise interactions between the elementary constituents. Those systems can be mapped onto two graphs $G_{1}$ and $G_{2}$. We assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices. For simplicity we focus here on undirected simple graphs, without loops. The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal. Additionally, we assume an isotropic structure for all two-body interactions, such that the valence of the graphs is constant and equal to $k$. Diestel (2017) In particular,
$\sum_{i=1}^{N_{G}}[A_{n}]_{ij^{\prime}}=\sum_{j=1}^{N_{G}}[A_{n}]_{i^{\prime}j}=k,\hskip 22.76219pt\forall i^{\prime},j^{\prime}=1,\cdots,{N_{G}},$ (II.1)
where $n=1,2$. We highlight that, as proven in recent work, Scholes (2023); Scholes and Amati (2024) the present formalism is robust even in the presence of disorder in the network, which may violate the condition of exact regularity. Consequently, the theory discussed in the paper is also applicable to quasi-regular resources, provided analytical results on matrix algebra are replaced by their more general numerical analogs. The direct sum of the two adjacency matrices is
@@
where we labeled by $\mathbb{0}_{N_{G}}\in M_{N_{G}}(\mathbb{N})$ the null matrix on the off-diagonal blocks. As proposed in Ref. Scholes, 2024, to define QL bits we switch on an interaction between the two subgraphs, and define the resource
@@
As discussed in Refs. Scholes, 2023, 2024, resources with structure as in ?? are particularly appealing for information processing. In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states. By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information processing that can be controlled by network manipulation. We will elaborate on this important concept further in the manuscript.
@@
where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$. The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ?? with eigenvalues $\{k-l,k+l\}$, respectively. The current arrow notation for the computational basis helps prevent any ambiguity with additional numeric subscripts and superscripts, which are used here and in the following sections for different purposes. In prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2024) ?? describes another basis choice, equally valid. The present approach establishes a direct connection between the computational basis and the emergent states of the composite graph described by ??. In the following we show that the mapping in ?? leads to a well-defined Hilbert space. First of all, it is straightforward to show that the $\psi_{\sigma}$’s are mutually orthogonal and normalized:
@@
In this section we summarized a list of prescriptions for defining a mapping from a bipartite graph to the Hilbert space of a single qubit, and related properties. Scholes and Amati (2024); Scholes (2024) In the next ?? we expand on an idea proposed in Ref. Scholes and Amati, 2024, to represent entanglement and superposition by coupling together multiple graph resources. Formal derivations on how to implement information processing and define a QL probabilistic model in this framework, follow in the later ????, respectively.
@@ II.2.1 Formalism
Various representations of the quantum states are possible with respect to classical network structures. For a small number of QL bits the representation defined in Ref. Scholes, 2024 is useful. In that work the Bell states, involving two QL bits, are designed to give insight on proof-of-principle examples of entanglement. In Ref. Scholes and Amati, 2024 we propose a formal representation of states applicable to an arbitrary number of QL bits. This is based on the Cartesian product of single-QL-bit graphs. Consistently with that work, we define here as a simple generalization of ?? the resource for a system of ${N_{\mathrm{QL}}}$ bits
@@
are eigenvectors of ??. Here, different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\cdots,\sigma^{(N_{\mathrm{QL}})}\right\}$, while the components of ?? are expressed via the multi-index $\bm{i}=\left\{i^{(1)},\cdots,i^{(N_{\mathrm{QL}})}\right\}$ as
@@
As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\{0,1\}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$. Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$. The action of $\mathcal{R}$ on $\psi_{\bm{\sigma}}$ is then given by
@@ II.2.2 Example: Bell states
$\psi^{(1)}_{[\uparrow\downarrow]}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{[\uparrow\downarrow]}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. The square brackets indicate that the sign on the right-hand side of each expression (+ or -) is chosen depending on the direction of the arrow on the left-hand side ($\uparrow$ or $\downarrow$, respectively). With ?? we can express
$\psi_{\uparrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\pm a_{2}\otimes b_{2}\end{pmatrix},$ (II.22a)
$\psi_{\downarrow,[\uparrow\downarrow]}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}&\pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}&\mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22b)
@@
$\Phi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
$\Psi^{\mathrm{Bell}}_{[\uparrow\downarrow]}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&a_{2}\otimes b_{2}\end{pmatrix},$ (II.24a)
@@
$\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix},$ (II.24b)
$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}&\mathbb{0}_{N}\\ \mathbb{0}_{N}&-a_{2}\otimes b_{2}\end{pmatrix},$ (II.24c)
$\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{N}&-a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}&\mathbb{0}_{N}\end{pmatrix}.$ (II.24d)
By applying the mapping from emergent states to the Hilbert space of QL bits, as discussed here, we can introduce the concept of QL gates for information processing. We propose an approach to implement these gates in the next ??.
III Quantum-like gates
In this section, we discuss the fundamental principles of QL information processing. Specifically, we demonstrate how the emergent states of QL resources can be transformed consistently to any arbitrary quantum gate. We then study how to translate these transformations into operations on the edges of the graph resources. This is particularly relevant for the future computational and experimental implementation of QL algorithms.

### v5

@@ Abstract
In previous work we have proposed a construction of quantum-like bits that could endow a large synchronizing classical system, for example of oscillators, with quantum-like function that is not compromised by decoherence. In the present paper we investigate further this platform of quantum-like states. Firstly, we discuss a general protocol on how to construct classical synchronizing networks that allow for emergent states. We then study how gates can be implemented on those states. This suggests the possibility of quantum-like information processing on a special class of many-body classical systems. Finally, we show that our approach allows for non-Kolmogorov interference, a feature that separates our model from a classical probabilistic system. This paper aims to explore the mathematical structure of quantum-like resources distilled from classical synchronizing systems, and shows how arbitrary gates can be implemented by manipulating many-body correlations.
@@ I Introduction
A large class of regular graphs encodes the correlations required to identify robust collective states. This general feature can be exploited for a variety of applications. For example, we discussed in Ref. Scholes, 2025a how this principle can support the engineering of molecular delocalization. The same reference, along with Refs. Krebs and Shaheen, 2011; Brouwer and Haemers, 2011, offers the interested reader background information on the mathematical formalism of regular and expander graphs, as well as their spectral properties. In the spirit of quantum-inspired computing, we showed in recent work that network correlations can serve as resources for quantum information processing. Scholes (2020, 2023, 2024); Scholes and Amati (2025) We focused, in particular, on many body systems whose correlations obey the graph structure shown in panel a) of ??. Here, two densely connected regions (in blue) of a network are mutually interacting via sparse links (in red). A spectral analysis of this class of graphs highlights the presence of two isolated emergent states. These an can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??). The robustness of emergent states against noise appears to be a direct consequence of synchronization. Scholes (2023, 2020) This intriguing feature is related to the occurrence of classical synchronization phenomena. Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) We showed in Ref. Scholes, 2024 that the emergent states of connected expander graphs are isomorphic to the Hilbert space of a qubit. This led us to introducing the notion of quantum-like (QL) bits within the context of classically-synchronizing networks. In Ref. Scholes and Amati, 2025 we built on these results and proposed a general framework to construct resources for multiple QL bits. In particular, we showed that the Cartesian product ($\square$) of ${N_{\mathrm{QL}}}$ bipartite graphs obeys the same dimensionality of a register of ${N_{\mathrm{QL}}}$ qubits [see panel b) of ??].
In the present paper, we build on the foundational work of Ref. Scholes, 2024; Scholes and Amati, 2025; Scholes, 2023, by proposing a self-contained formalization of the concepts of QL states and gates and their application to classical many-body systems. As sketched in ??, we present a framework to map the emergent states of a classical synchronizing network of oscillators onto a qubit Hilbert space. We highlight a connection between emergent states and steady-state synchronization dynamics, and we show that the latter is robust even in presence of stochastic disorder in the network. This makes the emergent states viable candidates for QL information processing, which we implement by transforming the matrix of the two-body correlations between the oscillators. Our analysis is general and applies to any classical system whose two-body correlations align with the network structures discussed later in this paper. However, to make the approach more concrete, we provide explicit examples in this manuscript by examining the synchronization dynamics of the Kuramoto model. This is a non-symplectic classical system consisting of a network of nonlinear oscillators, whose collective properties have been extensively explored in the literature. Kuramoto (1975); Kawamura et al. (2010); Scholes (2022); Rodrigues et al. (2016)
Our approach leads to a natural correspondence between “QL gates” transformations and the steady states of the classical oscillatory dynamics. Additionally, we propose a variety of experimental resources that can be utilized to test the formalism. This supports the applicability of the approach for information protocols involving low-dimensional Hilbert spaces, where the exponentially large complexity of the classical phase space can be handled by available experimental resources. Finally, we connect our perspective to the QL probabilistic framework developed by Khrennikov. Khrennikov (2007a, 2019, 2010) This theory formalizes the notion that genuine quantum systems are not required to process information in a QL way. Instead, to observe nonclassical interference it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and projective measurements.
@@
The present paper is structured as follows. In ?? we summarize our previous work describing how to construct a single-QL-bit Hilbert space. In ?? we discuss its generalization to higher dimensions, and we show explicitly in ?? how to construct Bell states for two QL bits. In ?? we show the fundamental role of emergent states in the synchronizing dynamics of the Kuramoto model as a paradigmatic classical system of synchronizing oscillators. In ?? we study the action of a complete set of QL gates, and we provide concrete examples of their action. In ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach. In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models. Khrennikov (2010) Finally, in ??, we propose a variety of experimental resources that could be utilized to implement and verify experimentally our formalism.
@@
In conjunction with Refs. Scholes, 2023; Scholes and Amati, 2025; Scholes, 2024, this work provides the mathematical foundation for a rigorous approach to QL information processing from classical many-body synchronizing systems.
@@
Figure 1: Illustration of the main concepts presented in this paper and in prior work by the authors. Scholes (2024); Scholes and Amati (2025) a) Many body systems consisting of networks of oscillators can easily exhibit classical synchronization phenomena. By structuring two-body correlations into two densely connected regions (blue) weakly interacting through sparse links (red), synchronization gives rise to pairs of emergent states. These states are well-distinguishable from a large band of random states at lower energies (dashed line in the sketch of the spectrum). In Ref. Scholes, 2024 we proposed a map from emergent states to the computational basis of a qubit. b) In Ref. Scholes and Amati, 2025, we demonstrated that the emergent states of the Cartesian product of a set of synchronizing graphs scale with the same dimensionality of a qubit system. In this paper, we expand over these concepts by bridging classical synchronization dynamics and the QL formalism. c) Additionally, we elaborate on the concept of QL information processing, and how this connects to steady-state synchronization dynamics. We show that operations on graphs, implemented through suitable transformations of their edges, are equivalent to gate transformations $\hat{V}_{\mathrm{g}}$ applied directly to the emergent states.
@@ II Quantum-like bits
In this section we derive a mapping from the emergent states of classical networks to the Hilbert space of interacting QL bits. While the concept has been already elucidated in Ref. Scholes, 2024, here we lay out a more formal description. The present analysis provides a complete description of the axioms of states and operators in quantum mechanics, within the notation of the emergent states. Later in the paper we show the direct connection between this abstract approach and the dynamics of many-body classical systems.
@@ II.1 Isolated system
Let us consider two classical many-body systems with pairwise interactions between the elementary constituents. The pair-wise interactions can be mapped onto two graphs $G_{1}$ and $G_{2}$. Here, we assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices. Also, we consider at this stage undirected simple graphs, without loops (while in the paper we will discuss graph transformations that break these symmetries). The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal. Additionally, we assume an isotropic structure for all two-body interactions, such that the valence of the graphs is constant and equal to $k$. Diestel (2017) In particular,
$\sum_{i=1}^{N_{G}}[A_{n}]_{ij^{\prime}}=\sum_{j=1}^{N_{G}}[A_{n}]_{i^{\prime}j}=k,\hskip 22.76219pt\forall i^{\prime},j^{\prime}=1,\ldots,{N_{G}},$ (II.1)
and $n=1,2$. A key property of these subgraphs is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs. Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) The direct sum of the two adjacency matrices is
@@
where we labeled by $\mathbb{0}_{N_{G}}\in M_{N_{G}}(\mathbb{N})$ the null matrix on the off-diagonal blocks. As proposed in Ref. Scholes, 2024, we define a QL bit resource by switching on an interaction between the two subgraphs, and define the resource
@@
As discussed in Refs. Scholes, 2023, 2024, resources with structure as in ?? are particularly appealing for information processing. In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states. By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information processing that can be controlled by network manipulation. We will elaborate on these important concepts further in the manuscript.
@@
where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$. As both the $A_{i}$’s and $C$ are exactly regular, then $a_{i}=(1,,1\ldots,1)^{\top}/\sqrt{2{N_{G}}}$ are vectors with constant weights. We keep, however, the notation general, to account for possible structural disorder in the network leading to this break of symmetry. The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ?? with eigenvalues $\{k-l,k+l\}$, respectively. Note that in prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2025) ?? describes another basis choice, equally valid. The present approach establishes a direct connection between the computational basis and the emergent states of the composite graph described by ??. In the following we show that the mapping in ?? leads to a well-defined Hilbert space. First of all, it is straightforward to show that the $\psi_{\sigma}$’s are mutually orthogonal and normalized:
@@
In this section we summarized a list of prescriptions for defining a mapping from a bipartite graph to the Hilbert space of a single qubit, and related properties. Scholes and Amati (2025); Scholes (2024) In the next ?? we expand on an idea proposed in Ref. Scholes and Amati, 2025, to represent entanglement and superposition by coupling together multiple graph resources. Formal derivations on how to implement information processing and define a QL probabilistic model in this framework, follow in the later ????, respectively.
@@ II.2.1 Formalism
Various representations of the quantum states are possible with respect to classical network structures. For a small number of QL bits the representation defined in Ref. Scholes, 2024 is useful. In that work the Bell states, involving two QL bits, are designed to give insight on proof-of-principle examples of entanglement. In Ref. Scholes and Amati, 2025 we propose a formal representation of states applicable to an arbitrary number of QL bits. This is based on the Cartesian product of single-QL-bit graphs. Consistently with that work, we define here as a simple generalization of ?? the resource for a system of ${N_{\mathrm{QL}}}$ bits
@@
are eigenvectors of ??. Here, different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\cdots,\sigma^{(N_{\mathrm{QL}})}\right\}\in\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}=\mathcal{S}^{N_{\mathrm{QL}}}$. The components of ?? are expressed via the multi-index $\bm{i}=\left\{i^{(1)},\cdots,i^{(N_{\mathrm{QL}})}\right\}$ as
@@
As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\mathcal{S}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$. Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$. The action of $\mathcal{R}$ on $\psi_{\bm{\sigma}}$ is then given by
@@ II.2.2 Example: Bell states
$\psi^{(1)}_{\uparrow/\downarrow}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\uparrow/\downarrow}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. The notation in ?? indicate that the upper arrow ($\uparrow$) on the left-hand side is associated to the positive sign on right-hand side, and similarly for $\downarrow$ and the negative sign. With ?? we can express
$\psi_{\uparrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \pm a_{2}\otimes b_{2}\end{pmatrix},\hskip 14.22636pt\psi_{\downarrow,\uparrow\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}\\ \mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22a)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
$\Psi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ a_{2}\otimes b_{2}\end{pmatrix},$ $\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix},$ (II.24a)
@@
$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ -a_{2}\otimes b_{2}\end{pmatrix},$ $\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ -a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix}.$ (II.24b)
So far, we have presented general results on the QL formalism and demonstrated that the approach enables an exact mapping to the Hilbert space of a qubit system. In the following ??, we explore how the abstract concepts of graph resources and emergent states relate to the classical dynamics of a system of oscillators. We discuss the resilience of this approach in the presence of noise in the network correlations and highlight the critical role of emergent states in the classical synchronizing dynamics.
III Classical synchronization
Here, we investigate classical synchronization in dynamical systems of oscillators, where two-body correlations are governed by a connectivity matrix with the structure described in ??. Systems that align with our formalism must, as a general requirement, involve classical many-body dynamics, where the two-body couplings between degrees of freedom are governed by a connectivity matrix $\mathcal{R}$, as outlined in ??. Additionally, the dynamics should guarantee that synchronization leads to a globally stable state that remains robust over long time scales. These features are satisfied by the non-symplectic dynamics of the Kuramoto model Kuramoto (1975); Kawamura et al. (2010); Scholes (2022); Rodrigues et al. (2016), which is characterized by its nonlinear equations of motion
$\dot{\theta}_{i}=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\sin(\theta_{j}-\theta_{i}),\hskip 19.91692pti=1,\cdots,2{N_{G}}.$ (III.1)
Here, the $\theta_{i}$’s are angular variables corresponding to $2{N_{G}}$ oscillators, defined on the vertices of a graph. The matrix $\mathcal{R}$ encodes the correlations between these degrees of freedom. The $\omega_{i}$’s denote the intrinsic oscillation frequencies. The time evolution of the angular variables of the model is shown in panel a) of ??. For our chosen system parameters (see caption of the figure) the system reaches a synchronized steady state, where all oscillators evolve at a uniform, constant phase. The synchronized steady state appears robust even in presence of disorder in the network, as shown in panel b) of the same figure. Here, white static disorder, with variance $\sigma=3$ and a zero mean, is added to all nonzero entries of the adjacency matrix $\mathcal{R}$, originally valued at one. Despite this strong perturbation, a collective synchronization state persists at long times, with only minor phase shifts affecting small subgroups of oscillators. This observation further supports the authors’ previous findings on the resilience of synchronization dynamics to noise–a property particularly appealing for robust information encoding and processing in the steady states of these classical dynamical systems. Scholes and Amati (2025); Scholes (2023, 2024)
To examine the role of the emergent states in the synchronization dynamics it is convenient to express the adjacency matrix through its spectral decomposition, as
$\mathcal{R}=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\lambda_{\bm{\sigma}}\left(\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}}^{\dagger}\right)+\sum_{i=1}^{M-2}\tilde{\lambda}_{i}\left(\tilde{\psi}_{i}\otimes\tilde{\psi}_{i}^{\dagger}\right),$ (III.2)
where $M=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of states. The first term in ?? represents the contribution of the emergent states $\psi_{\bm{\sigma}}$’s. The second term corresponds to the random part of the spectrum, with eigenvalues $\tilde{\lambda}_{i}$’s and corresponding eigenvectors $\tilde{\psi}_{i}$’s. Panel b) of ?? depicts the time evolution of the Kuramoto dynamics when only the random part of the spectrum is taken into account [that is, by imposing $\lambda_{\bm{\sigma}}\equiv 0$ $\forall\bm{\sigma}\in\bm{\mathcal{S}}$ in ??]. In this scenario, synchronization fails to emerge, with all degrees of freedom maintaining their initial random phases throughout the entire time evolution. This highlights that the structure of two-body correlations and emergent states, encoded in the adjacency matrix $\mathcal{R}$, play a critical role in achieving classical synchronization. In the next ??, we build on these idea by exploring the connection between dynamical synchronization and information processing. This approach provides a framework for linking the steady states of classical many-body systems to quantum information protocols.
Figure 5: a): Time evolution of a system with $602{N_{G}}=$ 60 oscillators governed by the Kuramoto model [??]. The angular variables $\theta_{i}$ are initialized uniformly in $[0,2\pi)$, and the intrinsic frequencies $\omega_{i}$ are sampled uniformly from $(0,1)$. The coupling constant is set to $\Gamma=10$, and the connectivity matrix $\mathcal{R}$ is structured as a regular graph, as described in ??, with the valencies of the diagonal and off-diagonal blocks set to $k=29$ and $l=15$, respectively. b): Time evolution of the system under the same parameters and initial conditions as in a), but with disorder introduced into the connectivity matrix $\mathcal{R}$. Specifically, each nonzero entry of $\mathcal{R}$ is perturbed by random white noise with zero mean and variance $\sigma=3$. c): Dynamics of the Kuramoto model with the connectivity matrix restricted to its random components [see discussion following ??]. The system parameters and initial conditions are the same as in a) and b).
IV Quantum-like gates
In this section, we explore the fundamental principles of QL information processing. In ??, we formalize how the emergent states of QL resources can be systematically transformed according to arbitrary quantum gates. We further investigate the correspondence between state transformations and network resources. In ??, we connect this abstract formalism to the dynamics of classical synchronization. Specifically, we demonstrate how QL gates can be practically realized in classical many-body systems of synchronizing oscillators and analyze the impact of these transformations on the long-time steady states of their dynamics.
IV.1 Formalism